# Yufei Bo, Shuo Shao, Meixia Tao

### v1

@@ Abstract
We consider multi-user semantic communications over broadcast channels. While most existing works consider that each receiver requires either the same or independent semantic information, this paper explores the scenario where the semantic information desired by different receivers is different but correlated. In particular, we investigate semantic communications over Gaussian broadcast channels where the transmitter has a common observable source but the receivers wish to recover hierarchical semantic information in adaptation to their channel conditions. Inspired by the capacity achieving property of superposition codes, we propose a deep learning based superposition coded modulation (DeepSCM) scheme. Specifically, the hierarchical semantic information is first extracted and encoded into basic and enhanced feature vectors. A linear minimum mean square error (LMMSE) decorrelator is then developed to obtain a refinement from the enhanced features that is uncorrelated with the basic features. Finally, the basic features and their refinement are superposed for broadcasting after probabilistic modulation. Experiments are conducted for two-receiver image semantic broadcasting with coarse and fine classification as hierarchical semantic tasks. DeepSCM outperforms the benchmarking coded-modulation scheme without a superposition structure, especially with large channel disparity and high order modulation. It also approaches the performance upperbound as if there were only one receiver.
@@ I Introduction
Leveraging the advantages of deep learning, semantic communication has mainly adopted neural networks (NNs) for semantic coding, which demonstrates superior performance in point-to-point communication scenarios. In particular, NN-based semantic coding has been used to substitute the conventional source coding and/or channel coding to transmit various source data types including speeches[6], texts[7, 8, 9], images[10, 11, 12, 13], videos[14, 15] as well as multi-modal data[16]. It also facilitates the execution of a wide range of intelligent semantic tasks, such as object detection[17], classification[10], and question-answering[16]. Different types of NNs have been employed for the semantic coding of different source data types. For example, the Resnet [18] is often used for image sources[10] while the Transformer[19] is often used for the text sources. Moreover, while earlier semantic communication systems often employ analog modulation to directly transmit through the channel the real-valued output of the NN-based semantic encoder[6, 10, 7], recent advancements have brought forth digital semantic communications by taking into account digital modulation explicitly in [20, 21, 22, 23, 24]. The works [20, 21] consider codebook-based quantization methods, and the work [22] designs a learned soft-to-hard quantizer. Moreover, the works [23, 24] introduce a novel joint coding-modulation framework for end-to-end design of digital semantic communication systems, which utilizes variational autoencoder (VAE) to output the transition probability from source data to discrete constellation symbols, and can approach the optimal probabilistic constellation shaping under Gaussian channels.
@@
Notation
Description
$\mathbf{S}_{1}$, $\mathbf{\hat{S}}_{1}$
Coarse-grained semantic information and its recovery at Receiver 1
$\mathbf{S}_{2}$, $\mathbf{\hat{S}}_{2}$
Fine-grained semantic information and its recovery at Receiver 2
$\mathbf{X}$, $\mathbf{\hat{X}}_{1}$, $\mathbf{\hat{X}}_{2}$
Observable image source and its recoveries respectively at Receiver 1 and Receiver 2
$\mathbf{U}_{1}$
Basic encoded feature vector
$\mathbf{U}_{2}$
Enhanced encoded feature vector
$\mathbf{R}$
Successive refinement vector
$\mathbf{Y}_{1}$, $\mathbf{Y}_{2}$
Inner constellation sequence and outer constellation sequence
$\mathbf{Y}$
The super-constellation sequence to be sent into the channel
$\mathbf{Z}_{i}$
Received sequence of Receiver $i$
$f_{\bm{\theta}_{1}}(\cdot)$
Basic semantic encoder
$f_{\bm{\theta}_{2}}(\cdot)$
Enhancement semantic encoder
$\mathbf{W}$, $\mathbf{b}$
Weight and bias of the LMMSE decorrelator
$\bm{\alpha}_{i}$
NN parameters of Modulator $i$
$g_{\bm{\eta}_{1}}(\cdot)$, $g_{\bm{\eta}_{2}}(\cdot)$
Semantic decoders for image recovery of Receiver 1 and Receiver 2
$g_{\bm{\psi}_{1}}(\cdot)$, $g_{\bm{\psi}_{2}}(\cdot)$
Semantic decoders for coarse classification of Receiver 1 and fine classification of Receiver 2
While substantial research has been devoted to point-to-point semantic communications[6, 7, 8, 9, 10, 11, 12, 14, 15, 23], recently there is a growing interest in semantic communications over multi-user channels. Many-to-one semantic communication systems with multiple transmitters and a single receiver are considered in the works [25, 26, 27, 16, 28, 29, 30]. Specifically, the work [25] introduces a distributed joint source-channel coding (JSCC) scheme for correlated image sources, where each source is transmitted through a dedicated and independent noisy channel to the common receiver. A distributed JSCC scheme for image transmission over Gaussian multiple access channel (MAC) is considered in [26]. The work [27] considers a multi-user fading channel and introduces a channel-transferable semantic communication for orthogonal frequency division multiplexing with non-orthogonal multiple access (OFDM-NOMA) system. Moreover, works [16, 28, 29] consider multi-user channels with a multi-antenna receiver, where [28] considers cooperative object identification, while [16] and [29] address the transmission of multimodal data. Additionally, a novel multiple access technology called model division multiple access (MDMA) is proposed in [30], where both uplink and downlink scenarios are considered.
@@
In this paper we investigate semantic communication over a two-user broadcast channel where the different but correlated semantic information needed by the two receivers exhibits a hierarchical structure. Specifically, the common observable source and the associated semantic information are modeled as a hierarchical structure. The receiver with poor channel condition wishes to recover the observable source and the coarse-grained semantic information while the receiver with good channel condition wishes to recover the observable source and the fine-grained semantic information. Our goal is to design an efficient digital semantic communication framework that can exploit such hierarchical structure of semantic information and accommodate the diverse channel conditions of different receivers.
@@
To this end, we propose a novel deep learning-based superposition coded modulation (DeepSCM) scheme for hierarchical semantic communications over Gaussian degraded broadcast channels. Inspired by the capacity achieving property of superposition codes for degraded broadcast channels[36], our DeepSCM scheme is able to extract uncorrelated features for different receivers from a common observable source and encode them into a superposition-structured constellation, so that each receiver can decode different levels of semantic information according to its own channel condition. Specifically, we first use two separate NN-based semantic encoders to encode the common observable source into a basic encoded feature vector for the poor receiver and an enhanced encoded feature vector for the good receiver respectively. Recognizing the existing correlation between them, we thus design a decorrelator. This decorrelator splits the enhanced encoded feature vector into two parts, one containing information highly correlated with the basic encoded feature vector, the other containing uncorrelated information. The latter is denoted as the successive refinement vector of the basic encoded feature vector, and is subsequently modulated for transmission. Moreover, we design a modulation strategy, associating the basic encoded feature vector with the inner layer of a superposition-structured constellation and the successive refinement vector with the outer layer. As a result, the poor receiver can successfully decode the inner layer and recover the first level of semantic information. At the same time, the good receiver can further decode the outer layer, therefore recovering the enhanced level of semantic information.
Figure 1: The framework of the proposed DeepSCM scheme.
(a) 4QAM×$\times$4QAM, $a=0.7$.
(b) 4QAM×$\times$4QAM, $a=0.8$.
(c) 4QAM×$\times$4QAM, $a=0.9$.
(d) 4QAM×$\times$16QAM, $a=0.7$.
(e) 4QAM×$\times$16QAM, $a=0.76$.
(f) 4QAM×$\times$16QAM, $a=0.9$.
Figure 2: Constellations after superposition with various values of $a$. The red diamonds indicate the inner constellation multiplied by the PAF, and the dots indicate the resulted super-constellation. All symbols are assumed to be uniformly distributed in this figure with $P=1$. Notably, (b) and (e) present rectangular 16QAM and 64QAM, respectively.
Our main contribution is to propose a novel NN-based architecture for hierarchical semantic communications, namely the DeepSCM scheme, according to the idea presented above. Our proposed scheme combines the benefits of NN-based coding and the classical superposition coding to enable efficient semantic communications over Gaussian degraded broadcast channels. To implement such a complex system, the following efforts have been made. First, to improve transmission efficiency, we design a novel module called the linear mininum mean square error (LMMSE) decorrelator to decorrelate the enhanced encoded feature vector with the basic encoded feature vector, obtaining the successive refinement vector of the basic encoded feature vector. Theoretical proof is further provided to establish that the successive refinement vector and the basic encoded feature vector can indeed be uncorrelated through optimization. Second, to ensure the convergence and the performance of the training process, we devise a three-stage training strategy for the training of the DeepSCM scheme according to the superposition structure. Extensive experiments on real-world datasets validate two performance advantages of our DeepSCM scheme. First, our proposed DeepSCM scheme can simultaneously approach the best achievable performances for both receivers, as if there were only one receiver. Second, it outperforms the coded modulation scheme without superposition structure, where the transmitter only uses a single semantic encoder that is trained jointly with the semantic decoders at the two receivers. This performance advantage grows as the channel signal-to-noise (SNR) gap between the two receivers widens and the modulation order increases.

### v2

@@ Abstract
We consider multi-user semantic communications over broadcast channels. While most existing works consider that each receiver requires either the same or independent semantic information, this paper explores the scenario where the semantic information desired by different receivers is different but correlated. In particular, we investigate semantic communications over Gaussian broadcast channels where the transmitter has a common observable source but the receivers wish to recover hierarchical semantic information in adaptation to their channel conditions. Inspired by the capacity achieving property of superposition coding, we propose a deep learning based superposition coded modulation (DeepSCM) scheme. Specifically, the hierarchical semantic information is first extracted and encoded into basic and enhanced feature vectors. A linear minimum mean square error (LMMSE) decorrelator is then developed to obtain a refinement from the enhanced features that is uncorrelated with the basic features. Finally, the basic features and their refinement are superposed for broadcasting after probabilistic modulation. Extensive experiments are conducted for two-receiver image semantic broadcasting with coarse and fine classification as hierarchical semantic tasks. DeepSCM outperforms the benchmarking coded-modulation scheme without a superposition structure as well as the classic separate source-channel coding baselines, especially with large channel disparity and high order modulation. It also approaches the performance upperbound as if there were only one receiver.
@@ I Introduction
Leveraging the advantages of deep learning, semantic communication has mainly adopted neural networks (NNs) for semantic coding, which demonstrates superior performance in point-to-point communication scenarios. In particular, NN-based semantic coding has been used to substitute the conventional source coding and/or channel coding to transmit various source data types including speeches[6], texts[7, 8, 9], images[10, 11, 12, 13], videos[14, 15] as well as multi-modal data[16]. It also facilitates the execution of a wide range of intelligent semantic tasks, such as object detection[17], classification[10], and question-answering[16]. Different types of NNs have been employed for the semantic coding of different source data types. For example, the Resnet [18] is often used for image sources[10] while Transformer[19] and BERT[20] is often used for text sources[7, 21]. Moreover, while earlier semantic communication systems often employ analog modulation to directly transmit through the channel the real-valued output of the neural semantic encoder[6, 10, 7], recent advancements have brought forth digital semantic communications by taking into account digital modulation explicitly in [22, 23, 24, 25, 26]. In particular, the works [25, 26] introduce a novel joint coding-modulation framework for end-to-end design of digital semantic communication systems. Therein, variational autoencoder (VAE) is applied to generate the transition probability from source data to discrete constellation symbols, which resolves the non-differentiable issue of digital modulation and learns the probabilistic shaping of constellations for the operating channel condition.
@@
Notation Description
$\mathbf{S}_{1}$, $\mathbf{\hat{S}}_{1}$ Coarse-grained semantic information and its recovery at Receiver 1
$\mathbf{S}_{2}$, $\mathbf{\hat{S}}_{2}$ Fine-grained semantic information and its recovery at Receiver 2
$\mathbf{X}$, $\mathbf{\hat{X}}_{1}$, $\mathbf{\hat{X}}_{2}$ Observable image source and its recoveries respectively at Receiver 1 and Receiver 2
$\mathbf{U}_{1}$ Basic EFV
$\mathbf{U}_{2}$ Enhanced EFV
$\mathbf{R}$ Successive refinement vector
$\mathbf{Y}_{1}$, $\mathbf{Y}_{2}$ Inner constellation sequence and outer constellation sequence
$\mathbf{Y}$ The super-constellation sequence to be sent into the channel
$\mathbf{Z}_{i}$ Received sequence of Receiver $i$
$f_{\bm{\theta}_{1}}(\cdot)$ Basic semantic encoder
$f_{\bm{\theta}_{2}}(\cdot)$ Enhancement semantic encoder
$\mathbf{W}$, $\mathbf{b}$ Weight and bias of the LMMSE decorrelator
$\bm{\rho}_{i}$ NN parameters of Modulator $i$
$g_{\bm{\eta}_{1}}(\cdot)$, $g_{\bm{\eta}_{2}}(\cdot)$ Semantic decoders for image recovery of Receiver 1 and Receiver 2
$g_{\bm{\psi}_{1}}(\cdot)$, $g_{\bm{\psi}_{2}}(\cdot)$ Semantic decoders for coarse classification of Receiver 1 and fine classification of Receiver 2
$n$ Number of channel uses
$\alpha$ Power allocation factor
While substantial research has been devoted to point-to-point semantic communications[6, 7, 8, 9, 10, 11, 12, 14, 15, 25], recently there is a growing interest in semantic communications over multi-user channels. Many-to-one semantic communication systems with multiple transmitters and a single receiver are considered in [27, 28, 29, 16, 30, 31, 32]. Specifically, the work [27] introduces a distributed joint source-channel coding (JSCC) scheme for correlated image sources, where each source is transmitted through a dedicated and independent noisy channel to the common receiver. A distributed JSCC scheme for image transmission over Gaussian multiple access channel (MAC) is considered in [28]. The work [29] considers a multi-user fading channel and introduces a channel-transferable semantic communication for orthogonal frequency division multiplexing with non-orthogonal multiple access (OFDM-NOMA) system. Moreover, the works [16, 30, 31] consider multi-user channels with a multi-antenna receiver, where [30] considers cooperative object identification, while [16] and [31] address the transmission of multimodal data. Additionally, a novel multiple access technology called model division multiple access (MDMA) is proposed in [32], where both uplink and downlink scenarios are considered.
@@
In this paper, we investigate semantic communication over broadcast channels where the different but correlated semantic information needed by multiple receivers exhibits a hierarchical structure. Specifically, the common observable source and the associated semantic information are modeled as a Markov chain. The receivers with poor channel conditions wish to recover the observable source and the coarse-grained semantic information while the receivers with better channel conditions wish to recover the observable source and the finer-grained semantic information. Our goal is to design an efficient digital semantic communication framework that can exploit such hierarchical structure of semantic information and accommodate the diverse channel conditions of different receivers.
To this end, we propose a novel deep learning-based superposition coded modulation (DeepSCM) scheme for hierarchical semantic communications over Gaussian degraded broadcast channels. The main idea is to utilize NNs to separately extract and encode the different levels of semantic information from the common source, then combine them using supposition coding before broadcasting. Thereby, each receiver can decode its desired level of semantic information from the superposition-structured symbols according to its own channel condition.
@@
Specifically, let us consider the two-user broadcast channel without loss of generality. We first use two separate neural semantic encoders to map the common observable source respectively into a basic encoded feature vector (EFV) (for the poor receiver) and an enhanced EFV (for the good receiver). Note that the two EFVs are generally correlated. As a classic capacity-achieving scheme for degraded broadcast channels[37], superposition coding, however, applies to independent data streams. We thus introduce a novel learning-based linear minimum mean square error (LMMSE) decorrelator module in our transmitter design. This decorrelator aims to extract the residual information from the enhanced EFV that is uncorrelated with the basic EFV, referred to as the successive refinement vector of the basic EFV. Theoretical proof is further provided to establish that the successive refinement vector and the basic EFV can indeed be uncorrelated through optimization. After that, we design a superposition modulation strategy, associating the basic EFV with the inner layer and the successive refinement vector with the outer layer. As a result, the poor receiver can decode the inner layer and recover the first level of semantic information. At the same time, the good receiver can further decode the outer layer, therefore recovering the enhanced level of semantic information. Moreover, to ensure the convergence and the performance of the training process, we devise a three-stage training strategy for the training of DeepSCM according to the superposition structure.
Figure 1: Framework of the proposed DeepSCM scheme.
Extensive experiments on real-world image datasets have validated the performance advantages of the proposed DeepSCM across various modulation orders, compression rates, and channel disparity levels. It outperforms the conventional coded modulation scheme without a superposition structure, where the transmitter uses a single semantic encoder that is jointly trained with the semantic decoders at the two receivers. This performance advantage grows as the channel signal-to-noise (SNR) gap between the two receivers widens and the modulation order increases. Furthermore, the proposed DeepSCM scheme can simultaneously approach the performance upperbound for both receivers, as if serving only one receiver.
@@
Throughout this paper, we use upper-case letters ($X$) and lower-case letters ($x$) to respectively denote random variables and their realizations. We use $h(X)$ to denote the differential entropy of the continuous variable $X$. The statistical expectation of $X$ is denoted as $\mathbf{E}[X]$. The covariance matrix of a vector random variable $\mathbf{P}$ is denoted as $\textit{Var}[\mathbf{P}]$, and $\textit{Cov}[\mathbf{P},\mathbf{Q}]$ denotes the cross-covariance matrix of $\mathbf{P}$ and $\mathbf{Q}$. We use $\mathcal{CN}(\mu,\sigma^{2})$ to denote the circularly symmetric complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. We use $\mathbf{I}_{k\times k}$ to denote the identity matrix with dimensions $k\times k$. We use $[M]$ to represent the set $\{1,...,M\}$ for $M\in\mathbb{Z}^{+}$, and occasionally use the short-hand Tx for a transmitter and Rx for a receiver. All important notations used in this paper are summarized in Table LABEL:notations.
II Overall Framework
This paper mainly focuses on semantic communication over two-user broadcast channels. The extension to $K$ ($K>2$) users shall be briefly discussed at the end of this section.
Fig. 1 illustrates the overall framework of the DeepSCM scheme over a two-user Gaussian degraded broadcast channel. Without loss of generality, we assume that Receiver 2 has a higher channel SNR than Receiver 1. There is an observable source $\mathbf{X}$ associated with implicit hierarchical semantic information, namely the coarse-grained semantic information $\mathbf{S}_{1}$ and the fine-grained semantic information $\mathbf{S}_{2}$. Naturally, they form a Markov chain as
$\mathbf{S}_{1}\to\mathbf{S}_{2}\to\mathbf{X}.$ (1)
Following the conventional setup of semantic communications[38, 12], each receiver needs to recover both the semantic information and the observable source, which reflects the practical demands in many real-world scenarios where both humans and machines are involved in decision-making. Receiver 1 requires the observable source $\mathbf{X}$ and the coarse-grained semantic information $\mathbf{S}_{1}$, the recovery of which we respectively denote as $\hat{\mathbf{X}}_{1}$ and $\hat{\mathbf{S}}_{1}$. Meanwhile, as Receiver 2 has a larger channel capacity, it requires the observable source $\mathbf{X}$ and the fine-grained semantic information $\mathbf{S}_{2}$, the recovery of which we respectively denote as $\hat{\mathbf{X}}_{2}$ and $\hat{\mathbf{S}}_{2}$.
Specifically, in this paper we focus on image semantic communications for both image recovery and classification. Namely, the observable source $\mathbf{X}\in\mathbb{R}^{d}$ is image data, where $d$ represents the dimension of the images. The coarse-grained semantic information $\mathbf{S}_{1}\in\{1,2,...,L_{1}\}$ is the label of coarse image classification with $L_{1}$ classes. Similarly, the fine-grained semantic information $\mathbf{S}_{2}\in\{1,2,...,L_{2}\}$ is the label of fine image classification with $L_{2}$ classes, where $L_{2}>L_{1}$. This hierarchical structure of the observable source and the semantics can well model the real-world scenarios. For instance, in wildlife monitoring, semantic information $\mathbf{S}_{1}$ can represent the general categories of animals, such as birds, bears, etc, while semantic information $\mathbf{S}_{2}$ can represent the specific species within those categories, such as sparrows, cardinals, black bears, brown bears, etc. Besides image classification, the above Markov-chain based hierarchical semantic information model (1) is also suitable for other tasks. For example, $\mathbf{S}_{1}$ could represent the location of prominent objects in an image relevant to the task of object detection, while $\mathbf{S}_{2}$ could indicate both the location and classification of the objects that are relevant to the task of image segmentation.
(a) 4QAM×$\times$4QAM, $\alpha=0.7$.
(b) 4QAM×$\times$4QAM, $\alpha=0.8$.
(c) 4QAM×$\times$4QAM, $\alpha=0.9$.
(d) 4QAM×$\times$16QAM, $\alpha=0.7$.
(e) 4QAM×$\times$16QAM, $\alpha=0.76$.
(f) 4QAM×$\times$16QAM, $\alpha=0.9$.
Figure 2: Constellations after superposition with various values of $\alpha$. The red diamonds indicate the inner constellation multiplied by the PAF, and the blue dots indicate the resulted super-constellation. All symbols are assumed to be uniformly distributed in this figure with $P=1$. Notably, (b) and (e) present rectangular 16QAM and 64QAM, respectively.
The transmitter extracts semantic features in a hierarchical way, corresponding to the hierarchical sources intended for the two receivers. At the transmitter, there are two joint coding-modulation (JCM) blocks responsible for generating two constellation sequences in a superposition structure. Specifically, the first JCM block generates the inner constellation sequence $\mathbf{Y}_{1}=(Y_{1}^{1},Y_{1}^{2},...,Y_{1}^{n})\in\mathbb{C}^{n}$, which is intended for both receivers and carries the first level of semantic features to recover the observable source $\mathbf{X}$ and the coarse-grained semantic information $\mathbf{S}_{1}$. Here, $n$ denotes the number of channel uses. The second JCM block generates the outer constellation sequence $\mathbf{Y}_{2}=(Y_{2}^{1},Y_{2}^{2},...,Y_{2}^{n})\in\mathbb{C}^{n}$, which is intended for Receiver 2. The outer constellation sequence carries the additional semantic features, which together with $\mathbf{Y}_{1}$, recovers the observable source $\mathbf{X}$ and the fine-grained semantic information $\mathbf{S}_{2}$.
The first JCM block consists of a basic semantic encoder and a modulator. The basic semantic encoder $f_{\bm{\theta}_{1}}$ with parameters $\bm{\theta}_{1}$ extracts and encodes the basic semantic features from $\mathbf{X}$, outputting a $2n$-dimensional real-valued basic EFV $\mathbf{U}_{1}\in\mathbb{R}^{2n}$
$\mathbf{U}_{1}=f_{\bm{\theta}_{1}}(\mathbf{X}).$ (2)
Then, a probabilistic modulator with parameters $\bm{\rho}_{1}$ generates $\mathbf{Y}_{1}$ from $\mathbf{U}_{1}$ by first learning a transition probability $p(\mathbf{y}_{1}|\mathbf{u}_{1},\bm{\rho}_{1})$ then randomly sampling a sequence according to this transition probability as in our previous work [25].
The second JCM block consists of an enhancement semantic encoder, an LMMSE decorrelator and a modulator. The enhancement semantic encoder $f_{\bm{\theta}_{2}}$ with parameters $\bm{\theta}_{2}$ generates an enhanced EFV $\mathbf{U}_{2}\in\mathbb{R}^{2n}$ from $\mathbf{X}$, which contains the semantic features of $\mathbf{X}$ and $\mathbf{S}_{2}$
$\mathbf{U}_{2}=f_{\bm{\theta}_{2}}(\mathbf{X}).$ (3)
Considering the redundancy in $\mathbf{U}_{2}$ due to the hierarchical semantics, the LMMSE decorrelator projects $\mathbf{U}_{2}$ onto the space of $\mathbf{U}_{1}$, resulting in the successive refinement vector $\mathbf{R}$ of $\mathbf{U}_{1}$, so that $\mathbf{R}$ and $\mathbf{U}_{1}$ can be uncorrelated. A detailed discussion on the LMMSE decorrelator will be further provided in Section III. Then, in the same manner that $\mathbf{Y}_{1}$ is generated, a probabilistic modulator with parameters $\bm{\rho}_{2}$ learns the transition probability $p(\mathbf{y}_{2}|\mathbf{r},\bm{\rho}_{2})$ then randomly samples $\mathbf{Y}_{2}$ based on $\mathbf{R}$.
Figure 3: NN architecture of the first JCM block.
In this paper, digital transmission is carried out with $M$-QAM modulation by superposing $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$. Each element in $\mathbf{Y}_{1}$ takes values from an $M_{1}$-QAM constellation $\mathcal{C}_{1}$, and each element in $\mathbf{Y}_{2}$ takes values from an $M_{2}$-QAM constellation $\mathcal{C}_{2}$, where $M_{1}$ and $M_{2}$ satisfy $M_{1}\times M_{2}=M$. The output of the transmitter, which we denote as $\mathbf{Y}\in\mathbb{C}^{n}$, is formed by
$\mathbf{Y}=\sqrt{\alpha}\mathbf{Y}_{1}+\sqrt{1-\alpha}\mathbf{Y}_{2},$ (4)
where $\alpha$ denotes the power allocation factor (PAF), and we set $\alpha\in(0.5,1)$ so that the inner constellation is truly “inner”. Note that before superposition, we respectively scale $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ so that they meet an average transmit power constraint $P$, i.e., $\frac{\|\mathbf{Y}_{1}\|^{2}}{n}=\frac{\|\mathbf{Y}_{2}\|^{2}}{n}=P$. The average transmit power constraint $P$ is again enforced after superposition
$\frac{\|\mathbf{Y}\|^{2}}{n}=P.$ (5)
Receiver 1 receives a sequence $\mathbf{Z}_{1}=\mathbf{Y}+\bm{\epsilon}_{1}$, where $\bm{\epsilon}_{1}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}_{1}\mathbf{I}_{n\times n})$. Similarly, Receiver 2 receives a sequence $\mathbf{Z}_{2}=\mathbf{Y}+\bm{\epsilon}_{2}$ with $\bm{\epsilon}_{2}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}_{2}\mathbf{I}_{n\times n})$ and $\sigma_{1}>\sigma_{2}$. The channel condition of Receiver $i$ for $i$ in $\{1,2\}$ is characterized by the channel SNR, which is defined as $\frac{P}{\sigma_{i}^{2}}$. Fig. 2 displays examples of superposed inner and outer constellations (which we henceforward call super-constellations). Fig. 2(a), 2(b), and 2(c) depict the super-constellation resulting from the superposition of two 4QAM constellations, denoted as the 4QAM×$\times$4QAM super-constellation. Fig. 2(d), 2(e), and 2(f) illustrate the super-constellation obtained by superposing a 4QAM inner constellation and a 16QAM outer constellation, denoted as the 4QAM×$\times$16QAM super-constellation. As a side note, the $M$-QAM digital modulation can be easily extended to other modulation schemes, such as $M$-PSK.
At each receiver, two neural decoders are deployed in parallel to recover $\mathbf{X}$ and $\mathbf{S}_{i}$, respectively, from $\mathbf{Z}_{i}$ ($i\in{1,2}$). That is,
$\mathbf{\hat{S}}_{i}$ $=g_{\bm{\psi}_{i}}(\mathbf{Z}_{i}),$ (6)
$\mathbf{\hat{X}}_{i}$ $=g_{\bm{\eta}_{i}}(\mathbf{Z}_{i}),$ (7)
where $\bm{\psi}_{i}$ parameterizes Receiver $i$’s decoder for classification and $\bm{\eta}_{i}$ parameterizes its decoder for image recovery.
The above DeepSCM framework can be easily extended to $K$ ($K>2$) users. More specifically, consider the observable source $\mathbf{X}$ with implicit hierarchical semantic information $\mathbf{S}_{1}$, $\mathbf{S}_{2}$, …, $\mathbf{S}_{K}$, forming the Markov chain
$\mathbf{S}_{1}\to\mathbf{S}_{2}\to...\to\mathbf{S}_{K}\to\mathbf{X}.$ (8)
Receiver $i$, for $i\in[K]$, requires the source data $\mathbf{X}$ and the semantic information $\mathbf{S}_{i}$. Similar to the two-user case, there are $K$ JCM blocks at the transmitter. For each JCM block $i$, for $i\in[K]$, it outputs the feature vector $\mathbf{U}_{i}\in\mathbb{R}^{2n}$, which contains the semantic features of $\mathbf{X}$, $\mathbf{S}_{1}$, …, $\mathbf{S}_{i}$. The feature vector $\mathbf{U}_{1}$ is directly modulated into the constellation sequence $\mathbf{Y}_{1}\in\mathbb{C}^{n}$, while for $i\geq 2$, $\mathbf{U}_{i}$ first goes through decorrelation to be projected as the successive refinement vector $\mathbf{R}_{i}$ of $\mathbf{U}_{i-1}$, which is then modulated to produce $\mathbf{Y}_{i}\in\mathbb{C}^{n}$. Each constellation sequence is normalized before they are superposed to form the output super-constellation sequence $\mathbf{Y}$
$\mathbf{Y}=\sum_{i=1}^{K}\sqrt{\alpha_{i}}\mathbf{Y}_{i},$ (9)
where $\alpha_{i}$ denotes the PAF for $\mathbf{Y}_{i}$, satisfying $\alpha_{1}>\alpha_{2}>...>\alpha_{K}$ and $\sum_{i=1}^{K}\alpha_{i}=1$. The superposed signal is normalized to the average power of $P$ before transmission. The receivers are similarly designed to those in the two-user case, with each receiver employing two parallel neural decoders to respectively recover the common source data and the desired-level of semantic information.
III Transmitter Design
The core of our DeepSCM scheme lies in the transmitter design, which aims at extracting uncorrelated semantic features and generating the inner and outer constellation sequences in the superposition structure. This section is dedicated to describing the components of the transmitter as well as the training algorithm. We first discuss the basic building block of the transmitter, namely the JCM block. Then, in Section III-B, we explain the technical detail of the aforementioned LMMSE decorrelator. Finally, in Section III-C, we present the training algorithm of the whole system.
III-A Joint Coding-Modulation
Extending our previous work[25, 26] to a broadcast setting, we use two JCM blocks to enable the digital modulation of the two EFVs. The NN architecture of the first JCM block is shown in Fig. 3. The basic semantic encoder includes multiple Resnet blocks[18] to map an image $\mathbf{X}$ of dimension $W\times H\times C$ into the basic EFV $\mathbf{U}_{1}$ of dimension $2n$, where $W$, $H$ and $C$ respectively denote the width, height and channel of the image, and we have $d=W\times H\times C$.
We then modulate $\mathbf{U}_{1}$ to obtain $\mathbf{Y}_{1}$. The modulation process is learned as a probabilistic model to avoid the inherent nondifferentiability problem of the digital modulation. A multi-layer perceptron (MLP) with a $\text{PReLu}(\cdot)$ activation function and a normalization layer outputs the transition probability, denoted as $p(\mathbf{y}_{1}|\mathbf{u}_{1},\bm{\rho}_{1})$. Since $\mathbf{Y}_{1}=(Y_{1}^{1},Y_{1}^{2},...,Y_{1}^{n})$ is discretely distributed with $M_{1}^{n}$ possible values, the transition probability is a discrete probability distribution with $M_{1}^{n}$ categories. To simplify the learning process, we model each element of $\mathbf{Y}_{1}$ to be conditionally independent, reducing the total number of probability categories to be learned to $n\cdot M_{1}$. That is, for each element of $\mathbf{Y}_{1}$, the MLP respectively outputs $M_{1}$ probabilities. Based on this probability distribution, the constellation symbol generator samples a constellation symbol for this element by using the Gumbel-Softmax method[39] as detailed below.
As a differentiable sampling technique, the Gumbel-Softmax method is used to reparameterize the discrete variable $\mathbf{Y}_{1}$ as a deterministic function of an independent random variable and the transition probability, thereby enabling the backpropagation through one sample of $\mathbf{Y}_{1}$. Specifically, in the forward pass, the sampling process of $Y_{1}^{j}$, where $j\in[n]$, from the transition probability $p(\mathbf{y}_{1}|\mathbf{u}_{1},\bm{\rho}_{1})$ can be expressed as
$y_{1}^{j}=\mathbf{c}^{T}\cdot\text{one-hot}\left(\underset{m\in[M_{1}]}{\mathrm{argmax}}[\tau_{jm}+\log q_{jm|\mathbf{u}_{1},\bm{\rho}_{1}}]\right),$ (10)